The 2010 American Community Survey estimates that 14.6% of Americans live below the poverty line, 20.7% speak a language other than English at home and 4.2% fall into both categories. Answer the following questions give all percents to the nearest whole percent) a) Are living below the poverty level and speaking a language other than English at home disjoint events? O Yes N。 b) Just in case you were guessing in part a), if the events were disjoint, what percent would fall into both categories? (Give your answer to the nearest whole percent) percant c)Are living below the poverty level and speakinga language other than English at home independent event? O Yes No d) Just in case you were guessing in part c), if the events were independent, what percent would fall into both categories? (Give your answer to the nearest whole percent) X percent e) If 1 choose tro Americans at random, what is the probability that both live below the poverty line? (Give your answer to the nearest whole percent) percent f) If 1 choose two Americans at random, what is the probability that both speak English at home? (sive your answer to the nearest whole percent) percent g) If I choose two Americans at random, what is the probability that one speaks English at home and the other does not? (Sive your answer to the nearest whole percent) parcant
Answer #1

d)

for independence :% fall in both =(0.146*0.207)*100 =3 percent

e)

P(both live below poverty line) =(0.146*0.146)*100 =2 percent

f)

P(both speak english at home) =((1-0.207)*(1-0.207))*100=63 percent

g)

probability=(0.207*(1-0.207))*100=16 percent